# MMRel: A Relation Understanding Dataset and Benchmark in the MLLM Era

### v1

@@ 3.3 Statistics of MMRel
Table 2: Statistics of each subset in MMRel.
Domain Real SDXL Dall-E Total
[…]
Table 3: Number of objects and categories of relations in MMRel.
Object Spatial Action Comparative
855 5 378 9
Tab. 2 shows the statistics of inter-object relation data in MMRel. Specifically, MMRel comprises over 15,000 question-answer pairs across 7 subsets, spanning 3 domains and 3 categories of relations. Concretely, approximately 68% of the data consists of real images, while the remaining 32% comprises synthetic images (i.e., SDXL [26] and Dall-E [27]). In terms of relation categories, MMRel includes a larger scale of spatial and action relations compared to comparative relations. Thanks to the open-vocabulary capability of GPT-4V [28], MMRel guarantees a diverse range of objects and action relations, as shown in Tab. 3. Regarding comparative relations, we manually re-label VG images, resulting in a subset that, though relatively small in scale with limited diversity, ensures accurate annotations. It is important to note that our annotations strictly describe the objects and their relations, without including any attributes, so that the evaluation can accurately reflect the relation understanding capabilities. For example, if we introduce attributes into the question “Is blue bird right to yellow bird?”, and the MLLMs answers differ from the ground-truth, we cannot ascertain whether the discrepancy stems from a misunderstanding of attributes (i.e., blue and yellow) or inter-object relation (i.e., right). Inspired by [44], we recognize that unusual data are crucial for evaluating and enhancing MLLM capabilities. We therefore create an adversarial relation subset comprising 710 challenging question-answer pairs. This subset is designed to rigorously assess the relation understanding capabilities of MLLMs.
In summary, MMRel is featured with four distinctive characteristics as illustrated in Fig. 3(e): (i) It comprises large-scale and diverse relation data that are collected from different sources; (ii) It comprises high-quality relation annotations and negative-choice data for action relations; (iii) It adopts an effective and efficient evaluation framework [29] that employs straightforward Yes/No instead of other complicated discriminative or generative evaluations; (iv) It includes an adversarial subset designed to evaluate the MLLMs’ ability to understand challenging relations.
4 Experiments
In this section, we systematically analyze the effectiveness of Multi-Modal Relation Understanding (MMRel). We first describe the experiment setups in Sec. 4.1. Subsequently, we employ MMRel to evaluate the relation understanding capabilities of multiple representative Multi-Modal Large Language Models (MLLMs) and hallucination mitigation techniques in Sec. 4.2. Finally, we employ MMRel to fine-tune LLaVA-1.5 [2] to examine how much the fine-tuning enhances their relation understanding capabilities, which is further discussed in Sec. 4.3.
Table 4: Evaluation results of existing MLLMs on our MMRel benchmark. Existing MLLMs can not handle the task of relation understanding well.
@@ 4 Experiments
4.1 Experiment Setups
MLLM Baselines. We conduct experiments with six representative MLLMs [2, 5, 7, 30, 31, 32]. Three of them, namely, LLaVA-1.5 [2], InstructBLIP [5], and Qwen-VL [7], are versatile and can work on a variety of vision-language tasks. However, they are trained with image-text pairs with coarse granularity [45, 46] and thus prone to generating hallucinations [47, 48]. We therefore select another three MLLMs that were designed for hallucination mitigation. As MMRel involves inter-object relations only, we select DOLA [30], OPERA [31], and VCD [32] that focus on mitigating object hallucinations and so are more suitable for evaluating their relation understanding capabilities. Notably, the techniques selected for managing object hallucinations do not involve additional training, ensuring fair comparisons.
Evaluation. For fair benchmarking, we evaluate all six MLLMs using their 7B versions. Specifically, LLaVA-1.5 [2] and InstructBLIP [5] incorporate Vicuna 7B as their language decoder [49], while Qwen-VL adopts Qwen-7B backbone [7]. Given the impressive performance of LLaVA-1.5 [2], we also employ DOLA [30], OPERA [31], and VCD [32] in our evaluation, all based on LLaVA-1.5 [2]. As MMRel adopts a Yes/No evaluation framework, we adopt the same metrics, accuracy, precision, recall, and F1-Score, as in POPE [29]. We follow all details of evaluation as VCD [32].
4.2 Evaluations with MMRel
We first employ the 15K question-answer pairs in MMRel to evaluate how MLLMs perform when facing three categories of relations from three data domains. As Tab. 4 shows, all six MLLMs face various challenges in relation understanding. Specifically, we observe that the performance on action relations is superior to that on spatial and comparative relations in real images. This disparity arises because the coarsely-aligned training data (all are real images) predominantly features action descriptions, whereas spatial and comparative terms are often represented by more vague expressions, such as near, next to, in front of, etc. Furthermore, because the synthetic images generated by SDXL [26] and DALL-E [27] have distributions that differ significantly from the pre-training data, their performance on action relations is lower than that observed with real images. Moreover, even techniques designed to address object hallucinations, such as those reported in [30, 31, 32], continue to struggle with relation understanding. Their limited performances underscore that the primary challenges in MMRel arise from understanding relations rather than recognizing objects. The performances in Tab. 4 could serve as baselines for further training-free methods.
Table 5: Statistics of training, test and adversarial sets in MMRel.
@@ 4.3 Fine-Tuning with MMRel
The large scale and high diversity of MMRel also enable it to fine-tune MLLMs to enhance their relation understanding capabilities. To verify this, we split the 15K data into training and test sets, as shown in Tab. 5. Here, we only adopt LLaVA-1.5 [2] as a strong baseline due to its popularity, outstanding performance, and versatility. To maintain the generality of the MLLMs and prevent overfitting to our training data, we have integrated the MMRel training set with 260K instruction data points from LLaVA [1]. Furthermore, adhering to the instruction template used in LLaVA [1], we have expanded our Yes/No answers into various formats to enhance the diversity of the dialogues (with more details in the Appendix). We strictly follow the steps for instruction tuning from LLaVA-1.5 [2]. The statistics of our adversarial subset are shown in Tab. 5.
Table 6: Fine-tuning LLaVA-1.5 with MMRel improves the performance significantly.
Real Spatial LLaVA-1.5 [2] 51.29 50.88 74.83 60.57
Finetune with MMRel 79.66 82.45 75.34 78.74
Action LLaVA-1.5 [2] 65.08 61.68 84.97 71.47
Finetune with MMRel 81.15 81.21 82.47 81.84
Comparative LLaVA-1.5 [2] 64.04 83.33 66.18 73.77
Finetune with MMRel 76.40 79.75 92.65 85.71
SDXL Spatial LLaVA-1.5 [2] 50.63 50.39 80.63 62.02
Finetune with MMRel 95.63 96.20 95.00 95.60
Comparative LLaVA-1.5 [2] 61.24 82.89 63.00 71.59
Finetune with MMRel 77.13 77.43 99.50 87.09
Dall-E Spatial LLaVA-1.5 [2] 65.46 51.57 89.60 65.46
Finetune with MMRel 94.06 94.95 93.07 94.00
Action LLaVA-1.5 [2] 57.69 57.94 66.83 62.07
Finetune with MMRel 71.79 67.04 89.60 76.69
Adversarial Mix LLaVA-1.5 [2] 50.35 47.10 70.50 56.47
Finetune with MMRel 67.52 60.79 81.37 69.59
Following instruction tuning with the MMRel training set, we observe significant and consistent improvements in performance across all three domains and relation categories, as detailed in Tab. 6. Additionally, such fine-tuning enhances the performance on the adversarial subset as well. To validate our hypothesis that integrating our MMRel with LLaVA instruction data [1] is beneficial to other perceptual tasks, we conduct comparative experiments using the MME benchmark [24] against the baseline LLaVA-1.5 model [2]. Experiments in Tab. 7 show that instruction tuning with MMRel achieves clear improvements for most of the MME tasks, verifying the scalability and applicability of such fine-tuning in various multi-modal tasks. The evaluation metrics of MME [24] are described in the Appendix.
Table 7: Fine-tuning LLaVA-1.5 with MMRel improves performances clearly on MME.
Method Position Existence Count Color Posters Celebrity Scene Landmark Artwork OCR Total Score
LLaVA-1.5 [2] 114.0 175.7 124.7 151.0 127.8 113.6 148.3 130.0 102.2 92.0 1279.3
FT 126.7 180.0 128.3 153.3 129.6 122.1 149.3 131.9 117.0 117.5 1355.7
@@ 5 Conclusion
We notice that Multi-Modal Large Language Models (MLLMs) encounter difficulties in understanding inter-object relations, a research area that remains under-explored. To address this gap, in this paper, we first provide a clear taxonomy and define relation hallucinations for inter-object relations, and further introduce the Multi-Modal Relation Understanding (MMRel) dataset. MMRel is specifically designed to probe inter-object relations in MLLMs and includes a challenging adversarial subset featuring highly unusual relations. Thanks to its large-scale, high-quality and diverse data, MMRel serves as a versatile tool for evaluating and fine-tuning MLLMs, enhancing their capabilities in relation understanding and other vision-language tasks. Extensive experiments across various popular MLLMs validate MMRel’s effectiveness.
Acknowledgments and Disclosure of Funding
This study is supported by the Interdisciplinary Graduate Programme, Nanyang Technological University. We thank students from Nanyang Technological University: Zhixiao Jin, Siwei Huang, Guanqiao Fu, Yu Wei, and Taorui Wang for establishing the MMRel dataset.
Appendix
Appendix A Limitations and Future Work
Despite conducting extensive experiments to verify the effectiveness of MMRel, our fine-tuning efforts are currently limited to using LLaVA-1.5 [2] as the base model. Besides, we only adopt Yes/No evaluation just like POPE [29]. Future work should consider incorporating generative evaluation metrics, such as those used in CHAIR [47], to assess object hallucinations more comprehensively. Furthermore, the dataset for comparative relations remains relatively limited, thus, there is a need to collect and annotate more data within this subset. Lastly, the taxonomy of inter-object relations could be expanded to include additional categories.
Appendix B MMRel Dataset
Sample Images. We specifically aim to generate images featuring unusual inter-object relations in multiple styles using DALL-E-3 [27]. Additional examples of synthetic images from Dall-E-3 [27] are shown in Fig. A1.
Figure A1: Illustration of sample images generated by Dall-E. These images contain unusual relations with different styles.
SemiDC via GPT-4V [28]. We employ an in-context learning paradigm to label images in the Visual Genome (VG) dataset [12] using GPT-4V [28]. The specific prompts used are illustrated in Fig. A2
Figure A2: In-context learning prompts for labeling images from VG [12].
SemiDC via Dall-E-3 [27]. We use the following prompts as inputs for DALL-E 3 [27] to generate images: a style photo depicting a strange or unusual object-object relation: inter-object relations. Here, style specifies the image style, selected from photo-realistic, watercolor, abstract, and oil painting, and inter-object relations describes the specific relation we aim to depict. For example, the prompt for generating the first image in Fig. A1 is: an abstract photo depicting a strange or unusual object-object relation: cat on toilet.
Appendix C Experiments
Data for Fine-tuning LLaVA-1.5 [2]. We integrate the training set of MMRel with 260K original instances from the total 665K dataset utilized for LLaVA [2] instruction tuning to maintain the versatility of MLLMs and prevent overfitting to our specialized training data. Specifically, this 260K subset includes question-answer instructions comprising 73K data points from COCO [50], 53K from VG [12], 72K from GQA [51], 22K from TextVQA [52], and 41K text-only data.
Question-answer Instructions. The original annotations are in the form of triplets, such as “cat on car”. We expand these into both discriminative and generative question-answer pairs.
The discriminative instructions are as follows:
Positive:
Q: Is there a cat on a car in the image? Please answer with one word.
A: Yes.
Negative:
Q: Is there a car on a cat in the image? Please answer with one word.
A: No.
The generative instructions are as follows:
Q: What is the spatial relation between a cat and a car in the image?
A: A cat is on a car.
Evaluation on MME [24]. MME [24] employs a Yes/No evaluation framework. Adhering to the original settings, we calculate the final score using the sum of accuracy and accuracy+, where accuracy is determined for each individual question and accuracy+ is assessed for each image, requiring correct answers to both associated questions. Accuracy+ serves as a stricter metric, offering a more comprehensive reflection of the MLLMs’ capabilities.

### v2

@@ 3.3 Statistics of MMRel
Tab. 2 shows the statistics of MMRel. Specifically, MMRel comprises around 22,500 question-answer pairs (15K Yes/No, and 7.5K Open-ended) across 7 subsets, spanning 3 domains and 3 categories of relations. Concretely, approximately 68% of the data consists of real images, while the remaining 32% comprises synthetic images (i.e., SDXL [46] and Dall-E [7]). In terms of relation categories, MMRel includes a larger scale of spatial and action relations compared to comparative relations. Thanks to the open-vocabulary capability of GPT-4V [1], MMRel guarantees a diverse range of objects and action relations, as shown in Tab. 2. Regarding comparative relations, we manually re-label VG images, resulting in a subset with accurate annotations. It is important to note that our annotations strictly describe the objects and their relations, without including any attributes, so that the evaluation can accurately reflect the relation understanding capabilities. For example, if we introduce attributes into the question “Is blue bird right to yellow bird?”, and the MLLMs answers differ from the ground-truth, we cannot ascertain whether the discrepancy stems from a misunderstanding of attributes (i.e., blue and yellow) or inter-object relation (i.e., right). Inspired by [5], we recognize that unusual data are crucial for evaluating and enhancing MLLM capabilities. We therefore create an adversarial relation subset comprising 710 challenging question-answer pairs. These pairs are meticulously manually selected from the complete MMRel to include images from all domains and encompass all relation categories. This subset is designed to rigorously assess the relation understanding capabilities of MLLMs, the statistics of which are described in Sec. 4.
Table 3: Evaluation results of various MLLMs on the MMRel. Results indicate that these MLLMs (including proprietary models and grounding-specific models) still struggle with inter-object relation understanding.
@@
GPT-4o [9] 81.07 78.13 88.29 82.90
Table 4: Statistics of training, test and adversarial (a challenging subset for testing) splits when MMRel serves as fine-tuning data, upper: number of images in each split, lower: number of distinct object and relation in each split.
@@
Domain Real SDXL Dall-E
Taxonomy Spatial Action Comparative Spatial Comparative Spatial Action
Property Object Relation Object Relation Object Relation Object Relation Object Relation Object Relation Object Relation
training split 475 5 111 304 149 9 75 4 40 2 164 5 151 141
test split 226 5 111 304 39 9 52 4 28 2 133 5 136 141
In summary, MMRel is featured with four distinctive characteristics as illustrated in Fig. 3(e): (i) It comprises large-scale and diverse relation data that are collected from different sources; (ii) It comprises high-quality relation annotations and negative-choice data for action relations; (iii) It adopts an effective and efficient evaluation framework [35] that employs straightforward discriminative Yes/No, and it provides a potential to evaluate generative open-ended questions; (iv) It includes an adversarial subset designed to assess the MLLMs’ ability to understand challenging relations.
4 Experiments
In this section, we systematically analyze the effectiveness of Multi-Modal Relation Understanding benchmark (MMRel). We first describe the experiment setups in Sec. 4.1. Subsequently, we employ MMRel to evaluate the relation understanding capabilities of multiple representative Multi-Modal Large Language Models (MLLMs) and hallucination mitigation techniques in Sec. 4.2. Finally, we employ MMRel to fine-tune LLaVA-1.5 [39] to examine how much the fine-tuning enhances their relation understanding capabilities, which is further discussed in Sec. 4.4.
4.1 Experiment Setups
MLLM Baselines. We evaluate MMRel over nine representative MLLMs [39, 18, 4, 12, 62, 17, 24, 29, 9]. Specifically, four MLLMs including LLaVA-1.5 [39], InstructBLIP [18], Qwen-VL [4], and MiniCPM-V [62] are open-sourced and capable of a variety of vision-language tasks. Nevertheless, they are trained with image-text pairs of coarse granularity [34, 32] and thus prone to generating hallucinations [50, 57]. Another three MLLMs were trained for hallucination mitigation. Since MMRel focuses on inter-object relations, we select DOLA [17], OPERA [24], and VCD [29] that aim for mitigating object hallucinations and are therefore suitable for evaluations on inter-object relations. In addition, we select another two MLLMs, namely, Shikra [12] that is trained for fine-grained grounding tasks and the latest proprietary GPT-4o [9] that demonstrates amazing capabilities on a bunch of vision-language tasks.
Evaluation. For benchmarking fairness, we evaluate all open-sourced MLLMs using their 7B (or similar scale) versions. Specifically, LLaVA-1.5 [39], InstructBLIP [18], and Shikra [12] incorporate Vicuna 7B as their language decoder [16]. MiniCPM-V [62] selects LLaMA-3 [20] and Qwen-VL adopts Qwen-7B backbone [4]. The three hallucination mitigation models including DOLA [17], OPERA [24], and VCD [29] are all based on LLaVA-1.5 [39]. For discriminative Yes/No evaluation framework, we adopt the same metrics, accuracy, precision, recall, and F1-Score, as in POPE [35]. For generative open-ended questions, our evaluations are LLM-assisted. We follow all details of evaluation as VCD [29].
4.2 MLLM Evaluations with MMRel
We employ all of the 15K Yes/No question-answer pairs in MMRel to evaluate how MLLMs perform while handling multimodal data with rich inter-object relations. As Tab. 3 shows, all nine MLLMs face various problems while handling relation understanding. Several points can be further drawn from the experimental results.
Regarding the types of inter-object relations, we can observe that MLLMs perform clearly better on action and comparative relations for real images. The better performance can be largely attributed to the coarsely-aligned MLLM training data (all are real images) that predominantly capture action and comparative descriptions instead of spatial terms which often appear with vague expressions such as near, next to, in front of, etc. From the perspective of synthetic vs. real data, we can see that MLLMs generally perform better on real images as compared with synthetic images by SDXL [46] (on comparative relations) and DALL-E [7] (on action relations). The better performance is largely attributed to the distribution gap between the two sets of synthetic images and the real images that were used in MLLM training.
Table 5: LLM-assisted open-ended evaluation of some MLLMs on our MMRel benchmark. The full mark is 10.
Domain Taxonomy Model Avg. Score
Real Spatial InstructBLIP [18] 1.27
LLaVA-1.5 [39] 1.31
Qwen-VL [4] 1.22
Action InstructBLIP [18] 5.28
LLaVA-1.5 [39] 5.26
Qwen-VL [4] 5.28
Comparative InstructBLIP [18] 5.05
LLaVA-1.5 [39] 5.12
Qwen-VL [4] 5.27
SDXL Spatial InstructBLIP [18] 1.23
LLaVA-1.5 [39] 1.20
Qwen-VL [4] 1.14
Comparative InstructBLIP [18] 5.50
LLaVA-1.5 [39] 5.05
Qwen-VL [4] 5.30
Dall-E Spatial InstructBLIP [18] 2.03
LLaVA-1.5 [39] 1.97
Qwen-VL [4] 1.96
Action InstructBLIP [18] 4.82
LLaVA-1.5 [39] 4.79
Qwen-VL [4] 4.80
With respect to open-sourced vs. proprietary MLLMs, we notice that GPT-4o [9] performs much better than all open-sourced MLLMs [39, 18, 4, 12, 62]. This demonstrates the critical role of large-scale data and large models in relation understanding, underscoring the great potential for further advancements in open-sourced MLLMs.
Regarding the recent works on hallucination mitigation, we observe that all three tested methods [17, 24, 29] remain problematic in relation understanding, indicating that the primary challenges in MMRel arise from relation understanding rather than object recognition.
Note the experiments in Tab. 3 can serve as baselines for future investigation on training-free methods.
4.3 LLM-assisted Open-ended Evaluation
Motivation. Even though discriminative Yes/No evaluation is widely adopted for its straightforward merit, evaluating generative open-ended answers is gaining traction because it more closely resembles natural conversations or dialogues. Thereafter, we additionally introduce an LLM-assisted open-ended evaluation for relation understanding, which constructs MMRel to be more comprehensive. More details about question-answer pairs are discussed in the supplementary materials.
Table 6: Fine-tuning LLaVA-1.5 with MMRel significantly improves relation understanding performance.
Domain Tax. Model Acc.(%) Prec.(%) Rec.(%) F1(%) #FP
Real Spat. LLaVA [39] 51.29 50.88 74.83 60.57 419
Finetune 79.66 82.45 75.34 78.74 93
Act. LLaVA [39] 65.08 61.68 84.97 71.47 274
Finetune 81.15 81.21 82.47 81.84 99
Comp. LLaVA [39] 64.04 83.33 66.18 73.77 9
Finetune 76.40 79.75 92.65 85.71 16
SDXL Spat. LLaVA [39] 50.63 50.39 80.63 62.02 127
Finetune 95.63 96.20 95.00 95.60 6
Comp. LLaVA [39] 61.24 82.89 63.00 71.59 26
Finetune 77.13 77.43 99.50 87.09 58
Dall-E Spat. LLaVA [39] 52.72 51.57 89.60 65.46 170
Finetune 94.06 94.95 93.07 94.00 10
Act. LLaVA [39] 57.69 57.94 66.83 62.07 98
Finetune 71.79 67.04 89.60 76.69 89
Adv. Mix LLaVA [39] 50.35 47.10 70.50 56.47 255
Finetune 67.52 60.79 81.37 69.59 169
LLM-assisted Evaluation. We adopt GPT-4o [9] to score the responses from MLLMs (range 0 to 10), which is similar to LLaVA-Bench [40]. The specific prompts are as in the supplementary materials.
Experiments and Analysis. As discussed in Sec. 4, methods for grounding tasks [12] and hallucination mitigation [29, 24, 17] still struggle with relation understanding, thus we evaluate the performances of three representative general-purpose MLLMs via open-ended questions. From the results in Tab. 5, we obtain the following conclusions: First, spatial relations are more challenging than action and comparative relation, which may caused by that training data describe coarse inter-object spatial relationships (i.e., near, behind, next to) instead of specific ones in MMRel (i.e., left, right, on, under). Second, action relations in the Dall-E domain pose more difficulties than action in the real domain because the data distribution of synthetic images differs from that of training data. Third, MLLMs perform not well on inter-object relation understanding. These observations are the same as those in Yes/No evaluations in Sec. 4, which demonstrate the effectiveness and fairness of the LLM-assisted open-ended evaluation. Nevertheless, the performance differences between MLLMs are relatively small, which cannot obviously reflect the comparative capabilities between them. We assume this limitation may come from (i) The LLM used for evaluation, GPT-4o [9], is not oracle for relation understanding, which is discussed in Sec. 4; (ii) LLM-assisted evaluations and open-ended questions are subjective and less straightforward compared to the clarity and simplicity of Yes/No evaluations. It is noteworthy that our proposed LLM-assisted evaluation provides a potential and reference for open-ended questions, which is better than only including Yes/No evaluations. We believe that open-ended questions are trends but are confined by the capabilities of current LLMs.
@@ 4.4 Fine-tuning with MMRel
Data Splits. The large-scale and diverse relation data in MMRel enable it to fine-tune MLLMs for enhanced relation understanding. We verify this by splitting the 15K Yes/No question-answer pairs into a training set and a test set that consists of 11K and 4K question-answer pairs, respectively. Table 4 shows the statistics of the split data. In addition, we also collect a subset of question-answer pairs of adversarial relations as discussed in Sec. 1 and 3. The subset was manually picked from data of different domains with different types of relations, aiming to gauge MLLMs’ relation understanding under challenging scenarios. The lower part of Tab. 4 shows this adversarial subset.
Setup. We adopt LLaVA-1.5 [39] as the baseline due to its popularity, outstanding performance, and versatility. While conducting the fine-tuning, we combine the MMRel training data with 260K instruction data from LLaVA-1.5 [40] to maintain the generality of the MLLMs and prevent overfitting to our training data. Following the instruction templates used in LLaVA [40], we expanded our Yes/No answers into various formats to enhance the diversity of the dialogues (more details are in the supplementary material). We strictly followed the instruction tuning steps from LLaVA-1.5 [39] as well.
Figure 5: Fine-tuning LLaVA-1.5 with MMRel improves performances clearly on MME, validating MMRel benefits other tasks.
Results. As Tab. 6 shows, fine-tuning with MMRel improves the capabilities of relation understanding significantly and consistently across all data domains and relation categories. In addition, fine-tuning improves the relation understanding of the adversarial subset as well. Specifically, the narrow definition of hallucinations refers to the scenario when MLLMs answer “yes” while the ground-truth is “no” or vice versa. The reductions in “False Positive (FP)” across most subsets in Tab. 6 indicate that fine-tuning with MMRel suppresses hallucinations effectively. Moreover, the fine-tuning with MMRel also mitigates the MLLMs’ bias toward answering “yes” rather than “no” [35, 38]. For instance, the “yes” ratio (for Real images) for spatial and action relations decreases to 45.68% and 52.28%, respectively, from the original 73.53% and 70.93% without the fine-tuning.
Discussion. To verify that fine-tuning MLLMs like LLaVA [40] with MMRel is beneficial to other perception tasks, we conduct experiments on an extensive benchmark MME [21] that is tailored to evaluate MLLMs from multiple perspectives. As Fig. 5 shows, the fine-tuning with MMRel improves most MME tasks as compared with the baseline LLaVA-1.5 [39]. This experiment verifies the scalability and applicability of the proposed MMRel while conducting fine-tuning in various multi-modal tasks. Please refer to supplementary material for more details regarding the evaluation metrics and experimental results on the MME benchmark [21].
@@ 5 Conclusion
We notice that Multi-Modal Large Language Models (MLLMs) encounter difficulties in understanding inter-object relations, a research area that remains under-explored. To address this gap, we contribute the Multi-Modal Relation Understanding benchmark (MMRel) with a clear taxonomy of inter-object relations and the definition of relation hallucinations. MMRel is specifically designed to probe MLLMs’ capabilities on relation understanding and includes a challenging adversarial subset featuring highly unusual relations. Thanks to its large-scale, high-quality, and diverse data, MMRel serves as a versatile tool for both evaluating and fine-tuning MLLMs, enhancing their capabilities in relation understanding and other perceptual vision-language tasks. Extensive experiments and corresponding analysis validate the effectiveness of MMRel.